OpenStudy (anonymous):

Two waves are given below, interferes each other what is the resultant wave and its amplitude?

OpenStudy (anonymous):

\[\psi 1 =20Cos(5x+25t) \] \[\psi 2 =20Cos(5x+25t+ \Pi/3)\]

OpenStudy (ujjwal):

Are the two waves travelling in the same direction?

OpenStudy (anonymous):

information given in the question buddy,,,, i guess Pi /3 is the angel

OpenStudy (ujjwal):

\(\pi\)/3 is phase difference.

OpenStudy (anonymous):

yeah same, phase difference between waves..

OpenStudy (anonymous):

resultant wave?

OpenStudy (ujjwal):

resultant wave = \(\psi =\psi 1 + \psi 2\) \[\psi=20\sqrt{3} \cos (5x +25 + \frac{\pi}{6})\]

OpenStudy (ujjwal):

so amplitude seems to be 20\(\sqrt3\)

OpenStudy (anonymous):

juss to add waves?

OpenStudy (anonymous):

is it?

OpenStudy (ujjwal):

yep

OpenStudy (anonymous):

 Just add these vectors to get the resultant amplitude.

OpenStudy (anonymous):

r u sure guys?

OpenStudy (anonymous):

Yep.